ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

11. Mathematical Physics

1991 MS Classification: 47, 81 Golovanova Nina F., Golovanov Anton A., The Moscow State Textile Academy, Russia Solution of the T-matrixs equation by the eikonal expansion method It is unsuccessfully to obtain by traditional methods the analytically solution of the main in quantum scattering theory T-matrixs equation T = V + VG_{0}T, if the interaction is strong. This equations solution that is grounded on the eikonal expansion and is valid for the wide variety in the potentials V and Green functions G_{0} will be presented. Some moments of that in the special case were published in Reference. Reference: N. F. Golovanova, V. Iskra, Description of elastic medium-energy proton-proton scattering in a wide range of angles, Phys. Lett. (1987) v. 187, n 1,2, 7-11.